The daily net profit for spaza shop follows a normal distribution with a mean of R220 and a standard deviation of R175.

a) The probability that the daily net profit exceeds R250 is (answer to 2 decimal places)   

b) The cut-off value for lowest 40 % of daily net profits is (in Rands and cents) 

 


1
Expert's answer
2022-05-24T08:32:03-0400

a)


"P(X>250)=1-P(Z\le\dfrac{250-220}{175})"

"\approx1-P(Z\le0.1714)\approx0.43"

b)


"P(X<x)=P(Z<\dfrac{x-220}{175})=0.4"

"\dfrac{x-220}{175}=-0.2533"

"x=220-175(0.2533)"

"x=175.67"